OpenStudy (anonymous):

(3a^4b)^2(-2a^4b)^4 Simplify

OpenStudy (anonymous):

(3a^4b)^2(-2a^4b)^4\[\left(3 a^4 b\right){}^{\wedge}2\left(-2 a^4 b\right){}^{\wedge}4 \]\[\left(9 a^8 b^2\right)\left(16 a^{16} b^4\right) \]\[144 a^{24} b^6 \]